OpenStudy (anonymous):

can u double integral e^y/x dxdy with limpets 4 to 0 and x^2 to 0

OpenStudy (anonymous):

\[\int\limits_{0}^{4}\int\limits_{0}^{x^2} e^{y/x} dx dy \]

OpenStudy (rational):

changing the order may help

OpenStudy (anonymous):

can u try and show me @rational

OpenStudy (anonymous):

Your double integral is not a valid region. You must have written it incorrectly. Look what happens when you integrate the exponential function. You will substitute a function of \(x\) into the antiderivative of the exponential, but then you have to integrate with respect to \(y\).

OpenStudy (anonymous):

thats what i have but lets say if u integral respect to y first then to x is it will be work ??

OpenStudy (anonymous):

Yes, integrating with respect to y first, and then x yields a finite result. You wouldn't even be able to integrate e^(y/x) with respect to x.

OpenStudy (anonymous):

okay so can u show me what is the integral of e^(y/x) respect to y

OpenStudy (rational):

you need to change the bounds also

OpenStudy (rational):

start by sketching the region for area of integration

OpenStudy (rational):

sketch below curves : y = 0 y = 4 x = 0 x = y^2 ???? (check this bound again)

OpenStudy (anonymous):

no no its x^2

OpenStudy (rational):

your bounds and the differential are not agreeing x = x^2 as a function is nonsense

OpenStudy (rational):

\[\large \int\limits_{0}^{4}\int\limits_{\color{red}{0}}^{\color{Red}{x^2}} e^{y/x} \color{Red}{dx} dy\]

OpenStudy (rational):

the bounds must show x as function of y since the differential is dx

OpenStudy (anonymous):

\[\int\limits_{0}^{4}\int\limits_{0}^{x^2} e^{y/x} dy dx\]

OpenStudy (anonymous):

how about this ??

OpenStudy (rational):

that looks fine

OpenStudy (rational):

\[\large \int\limits_{0}^{4}\int\limits_{\color{red}{0}}^{\color{Red}{x^2}} e^{y/x} \color{Red}{dy} dx\] like this ?

OpenStudy (rational):

\[\large \int\limits_{0}^{4}\left(\int\limits_{\color{red}{0}}^{\color{Red}{x^2}} e^{y/x} \color{Red}{dy}\right) dx\]

OpenStudy (rational):

first integrate the integral inside parenthesis treat "x" as constant

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

show me plz @rational i wanna know whats the integral of e^y/x

OpenStudy (rational):

x is a constant so it not much exciting as you are expecting

OpenStudy (rational):

\[\large \int\limits_{0}^{4}\left(\int\limits_{\color{red}{0}}^{\color{Red}{x^2}} e^{y/x} \color{Red}{dy}\right) dx =\int\limits_{0}^{4}\left(\dfrac{e^{y/x}}{1/x} \Bigg|_{\color{red}{0}}^{\color{Red}{x^2}} \right) dx\]

OpenStudy (anonymous):

than ??

OpenStudy (rational):

evaluate the bounds and you will be left with a single integral

OpenStudy (anonymous):

can u complete it

OpenStudy (rational):

\[\large \begin{align}\\ \int\limits_{0}^{4}\left(\int\limits_{\color{red}{0}}^{\color{Red}{x^2}} e^{y/x} \color{Red}{dy}\right) dx &=\int\limits_{0}^{4}\left(\dfrac{e^{y/x}}{1/x} \Bigg|_{\color{red}{0}}^{\color{Red}{x^2}} \right) dx \\~\\ &=\int\limits_{0}^{4}\left(x[e^{\color{Red}{x^2}/x} - e^{\color{red}{0}}] \right) dx \\~\\&= \int\limits_{0}^{4}\left(x[e^{x} - 1] \right) dx\end{align}\]

OpenStudy (anonymous):

okay than

OpenStudy (rational):

thats a single integral evaluate it

OpenStudy (anonymous):

okay just do it i wanna make sure looolz

OpenStudy (rational):

you should end up with that answer after evaluating the outer integral